OpenStudy (anonymous):

Help. x^2=(6)^2 + (3)^2 Using pythagorean theorem and square root

OpenStudy (anonymous):

\[x^2=36+9\] \[x^2=45\] \[x=\sqrt{45}=3\sqrt{5}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw